OpenStudy (anonymous):

A helmet was discounted 30%. If the regular price of the helmet is $80 what is the new price?

OpenStudy (anonymous):

The way I learned this was "Is over of, percent over 100" or "is/of = %/100 .. So the original price is 80, thats our of. and then we place 30 over 100. Cross multiplied and divide. So (80 x 30)/100 = your answer

OpenStudy (anonymous):

56

OpenStudy (anonymous):

or see 1 as 100%, then 30%=0,3. 1-0,3= 0,7 80*0,7= your answer

OpenStudy (anonymous):

take your result and subtract it from 80 you get 56

OpenStudy (anonymous):

yes